OpenStudy (anonymous):

All work for 5y -2y = 3y + 2

OpenStudy (tkhunny):

Well, do it. It's first a subtraction problem. Then it's a thought problem.

OpenStudy (141823):

I'll show you how to solve these. First, you want to simplify the problem, or combine like terms. 5y - 2y = 3y + 2 5y - 2y = 3y, so we can rewrite it as 3y = 3y + 2 You can already tell that there is a problem, because the statement is not true. It would have to be 3y = 3y, but theres the extra + 2 So, there is "No Solution" to the problem.

OpenStudy (anonymous):

first combine like terms. (5y-2y-3y) 0y=2 there is "no solution"